OpenStudy (anonymous):

what is the simplified form of the expression sqrt[6]{v}-sqrt{4v ^{2}} - sqrt{36v} + sqrt{v ^{2}}

OpenStudy (anonymous):

\[\sqrt[6]{v} - \sqrt{4v ^{2}}- \sqrt{36v} + \sqrt{v ^{2}}\]

OpenStudy (dan815):

\[\Large\sqrt[6]{v} - \sqrt{4v ^{2}}- \sqrt{36v} + \sqrt{v ^{2}}\]

OpenStudy (dan815):

simplify what u can first

OpenStudy (dan815):

is that maybe 6*sqrt(v)

OpenStudy (dan815):